OpenStudy (anonymous):

At a conference, the organizer determined that 32% of the attendees were between the ages of 18 and 29. Since there were 216 attendees in this age range, there were total attendees at the conference.

OpenStudy (thatonegirl_):

well 216 is 32% of some number we dont know. Correct?

OpenStudy (anonymous):

it is 69.12 i think

OpenStudy (thatonegirl_):

That doesn't even make sense if you think about it. The number is going to have to be higher because 216 is only 32% @mathkid911

OpenStudy (anonymous):

it is correct

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i see my internet was lagging

OpenStudy (thatonegirl_):

just remember that the word is is = and of is multiply. so from the word equation 216=.32x (put the percent into a decimal). Solve for x by dividing 216 by .32 and you get your answer.

OpenStudy (anonymous):

So 32% of what is 216... change percent to decimal : 32% = 32/100 = 0.32 " of " means multiply " what number " we will call x " is " means = 0.32 (x) = 216 -- divide both sides by 0.32 0.32x/0.32 = 216/0.32 x = 675 check... 32% of 675 = 216 0.32(675) = 216 216 = 216 (correct)

OpenStudy (anonymous):

it is 675

OpenStudy (anonymous):

if it is correct can i get a medal and get fanned?

OpenStudy (thatonegirl_):

You just copied what i wrote xD

OpenStudy (anonymous):

xD

OpenStudy (thatonegirl_):

lol

OpenStudy (anonymous):

I am smart enough to know you found that question from facebook xD